OpenStudy (anonymous):

A block of mass m slides down on rough inclined plane of angle x with the horizontal with an acceleration "a" along the plane the value of reaction is?

OpenStudy (anonymous):

@sami-21 @vikrantg4

OpenStudy (anonymous):

Normal reaction ?

OpenStudy (anonymous):

i think so...

OpenStudy (anonymous):

it's mg cos x :D

OpenStudy (anonymous):

but the answer should be

OpenStudy (anonymous):

\[m \sqrt{g^2 +a^2 - 2agsinx}\]

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

Is there any diagram in your book ?

OpenStudy (anonymous):

No

OpenStudy (vincent-lyon.fr):

Let us assume the block is moving down the plane. Normal reaction is: N = mg cos θ Tangential reaction (friction) is: T = mg sin θ - ma Hence total reaction is: R = √(N² + T²) = \(m \sqrt{g^2 +a^2 - 2ag\sin \theta}\)